# Learning Source Biases: Multi-sourced Misspecifications and Consequences

### v2

@@ 4.3 Unknown covariance matrix
[…]
Let $(\hat{b},\widehat{\Sigma})$ denote the KL minimizer. Define the DM’s misperception of source biases, as well as the distortion of his long-run behavior, as before. The next theorem pinpoints the difference between these quantities and their analogs in the baseline model where the true covariance matrix is known to the DM from the outset.
Theorem 2.
Let everything be as above. Then
  1. (i)
    $\Delta$ is the same as in the case where the DM knows $\Sigma^{*}$ from the outset, whereas $\widehat{\Sigma}=\Sigma^{*}+\Delta\Delta^{\top}$;
  2. (ii)
    $\mathbb{E}_{(b^{*},\Sigma^{*})}[\widehat{\emph{\text{BLP}}}(X)-\emph{\text{BLP}}^{*}(X)]$ equals $\delta$ times its counterpart in the case where $\Sigma^{*}$ is known from the outset, where $\delta\coloneqq(1+\Delta^{\top}\Sigma^{*-1}\Delta)^{-1}\in(0,1)$.
  3. (iii)
[…]
Part (i) of Theorem 2, which characterizes the DM’s long-run belief, adapts Theorem A.1 Case (III) of HKS to the current setting where the covariance matrix of the signals takes the form of $\Omega+\mathbf{1}_{N}\mathbf{1}_{N}^{\top}$. The remaining parts of the theorem are new. On the one hand, misspecified learning of the covariance matrix does not interfere with that of source biases and leaves the DM’s long-run misperceptions about the latter unaffected. On the other hand, misspecified learning of source biases makes the signals appear noisier to the DM than they truly are, in that the perceived covariance matrix equals its true counterpart, plus a rank-one update $\Delta\Delta^{\top}$ that is positive definite. [^10] The perceived noisiness of the signals as measured by, e.g., the determinant of their covariance matrix, is larger with compound misspecified learning than without, i.e., $\text{det}(\widehat{\Sigma})=\text{det}(\Sigma^{*})(1+\Delta^{\top}\Sigma^{*-1}\Delta)>\text{det}(\Sigma^{*})$, where the equality follows from Ding and Zhou (2007). Consequently, when making predictions about the true state, the DM relies less on signal realizations compared to when he knows the true covariance matrix from the outset. This is shown to mitigate the impact of misspecified learning on his behavior, resulting in a less biased prediction of the state on average.
To better understand the bias mitigation effect, we decompose the distortion of the DM’s long-run behavior as follows:
Effect (I) in the above expression is due to the misspecified learning of source biases and is shown to equal $-\overline{\Delta}$ in the baseline model. Effect (II) arises from the misspecified learning of the covariance matrix and has zero mean under the true signal distribution. Effect (III) captures the compound effect of misspecified learning. It equals
[…]
Turning to comparative statics: In light of Theorem 1(ii), it is unsurprising that increasing signal-to-noise ratios could have unintended consequences on the outcomes of misspecified learning. What is new here is the ambiguity associated with correcting misspecifications, whose comparative statics in the baseline model were straightforward. The ambiguity arises from compound misspecified learning: By correcting the misspecifications associated with familiar sources, we make the environment appear less noisy to the DM than before. This reduces the bias mitigation effect discussed above, i.e., $\delta=(1+\Delta^{\top}\Sigma^{*-1}\Delta)^{-1}$ increases, and in some cases could undo the benefit of misspecification corrections.
Theorem 2 sheds light on the behavioral and welfare consequences of the debiasing programs introduced toward the end of Section 3.2. Years of economic and psychological research have yielded mixed insights into the efficacy of these programs. For example, while the success of using one’s own mental resources to intercept implicit biases has been documented in prominent case studies (recall Footnote 8), recent meta-analysis of organizations that have implemented programs of similar kinds reveals a mixed pattern, leading Greenwald and Lai (2020) to conclude that “Convincing evidence for the effectiveness of these strategies is not yet available in peer-reviewed publications.” Even worse, efforts to suppress stereotypes or to reduce intergroup animosity may backfire (Kulik et al., 2000; Kalev et al., 2006), leading scholars including Bohnet (2016) to advocate alternative strategies such as evidence-based continuing learning. We take no stand on this debate and instead caution that even if misspecifications can be partially corrected (bring $\Delta_{i}$, $i\in\mathcal{M}$, closer to zero), their behavioral and welfare consequences may remain ambiguous due to the compound learning effect.
5 Further applications
In this section, we utilize our theoretical framework to analyze two further applications: misperceived media bias and project investment with misspecified learning.
5.1 Misperceived media bias
A Republican voter consults several news sources on an evolving issue. Some sources, such as the voter’s personal source and traditional news outlets, carry biases that he feels familiar with. Others, such as social media sites operated by unknown people or bots, are unfamiliar. Initially, the voter is uncertain about the horizontal biases of the unfamiliar sources, while holding dogmatic, wrong beliefs about the biases of some familiar sources. As someone who grew up watching Fox News, he believes that this source is more neutral than it actually is due to familiarity bias (Pennycook and Rand, 2021). He underestimates the bias associated with his own opinions and judgments, in part due to the bias blind spot and third-person effect (Pronin et al., 2002; Davison, 1983), and in part because he is working toward self-enhancement and would like to view himself as middle-of-the-road (Dennis, 1988). [^11] Bias blind spot refers to the tendency for people to recognize biases in human judgment except when that bias is their own. The third-person effect predicts that people perceive mass media messages to have a greater effect on others than on themselves. Finally, he believes in a liberal media bias like many other Republicans and displays excessive hostility towards CNN (Hassell et al., 2020). These cognitive biases are known to enhance with age, be resistant to corrective information, or even pass down through generations (Pronin et al., 2002; Jennings et al., 2009). They can be captured by letting ={personal source,Fox $\mathcal{M}=\{\text{personal source},\text{Fox News},\text{CNN}\}$ and $\tilde{b}_{i}<b_{i}^{*}$ $\forall i\in\mathcal{M}$ in our framework.
In each subsequent period $t=1,2,\cdots$, the voter takes a position $a_{t}\in\mathbb{R}$ and incurs a loss $-(a_{t}-\omega_{t})^{2}$ equal to the quadratic distance between this position and the true state of the world. The optimal position in the long run equals $\text{BLP}^{*}(X)$ under the correctly specified model and $\widehat{\text{BLP}}(X)$ under the misspecified model. The expected utilities generated by these positions, evaluated under the true signal distribution, equal $\text{EU}^{*}\coloneqq-\mathbb{E}_{(b^{*},\Sigma^{*})}\left(\omega-\text{BLP}^{*}(X)\right)^{2}$ and $\widehat{\text{EU}}\coloneqq-\mathbb{E}_{(b^{*},\Sigma^{*})}(\omega-\widehat{\text{BLP}}(X))^{2}$, respectively. The difference between them captures the welfare loss from misspecified learning.
Proposition 4.
$\widehat{\text{EU}}-\text{EU}^{*}$ equals $-\overline{\Delta}^{2}$ in the baseline model and $-\delta\overline{\Delta}^{2}$ in the environment described in Section 4.3.
Proposition 4, together with Theorems 1 and 2, fully describes the distortionary effects of misspecified learning on the voter’s long-run beliefs, behavior, and welfare. Since $\Delta_{i}=\tilde{b}_{i}-b_{i}^{*}<0$ $\forall i\in\mathcal{M}$ in the current example, $\overline{\Delta}<0$. The distortion of the voter’s worldview due to misspecifications constitutes a liberal media bias of magnitude $|\overline{\Delta}|$. Our results establish a mechanism through which this distortion can spill into the online news sphere and facilitate analysis of its behavioral and welfare consequences.
Hostile media effect and false polarization.
The hostile media effect refers to the phenomenon that opposing partisans perceive news coverage to be more biased against their side than it actually is. Evidence supporting the existence of this phenomenon abounds in political science, communication, and psychology (Feldman, 2014), [^12] In a classical experiment conducted by Arpan and Raney (2003), student subjects read a balanced story about their hometown college football team in one of three newspapers: home town, rival town, or neutral town outside the state. Subjects viewed the coverage as favoring the rival university, even if it was believed to be from the neutral town. More recently, Hassell et al. (2020) present journalists with balanced news stories that differ only in their ideological content (e.g., the party affiliation of the candidate running for office). Journalists exhibit no gatekeeping bias when selecting news stories to cover, despite the fact that many of them are ideologically liberal, thus debunking the “liberal media bias” that is widely perceived among Republicans. although research on learning and belief formation about new online sources has remained sparse. We demonstrate how the hostile media effect in traditional media can spill into the digital news sphere, as the voter in our model overperceives the bias of any liberal source while underperceiving the bias of any conservative source in the long run.
False polarization is defined as the difference between the perceived distance between the positions of two groups and the actual distance between them. Its presence in the American public was rigorously established by Levendusky and Malhotra (2016b). [^13] These authors used the survey data gleaned from a nationally representative sample that asked about the subjects’ self-placements on an issue scale, as well as their placements of a typical Democratic and Republican voter on that scale. The data revealed that Americans significantly misperceive the public to be more divided along partisan lines than they are in reality. To the extent that individual human beings can now express their opinions online at almost no cost, our model predicts a false polarization in the digital news sphere, as well as a causal relationship between misspecified news bias and false polarization. The latter is consistent with the literature’s common attribution of false polarization to partisan media exposure (Levendusky and Malhotra, 2016a).
Fact-checking.
DOOM stipulates that sources with initially unknown biases are irrelevant to the distortionary effect of misspecified learning. As we increase the signal-to-noise ratios of these sources, the voter’s expected utility under the true model and, hence, the misspecified model, increases unambiguously ($\text{EU}^{*}=-(1+\sum_{i\in\mathcal{S}}\nu_{i})^{-1}$ and $\widehat{\text{EU}}=\text{EU}^{*}-\overline{\Delta}^{2}$). The same cannot be said about sources with misspecified biases, as increasing their signal-to-noise ratios generally has an ambiguous effect on the welfare distortion $\overline{\Delta}^{2}$ from misspecified learning.
Our finding suggests that policy interventions that target misinformation and fake news through improving source accuracy (e.g., fact-checking) may entail unintended welfare consequences, especially when they operate by modulating sources that people already hold dogmatic misperceptions about. When it comes to lesser-known sources, fact-checking always improves welfare. To the extent that professional fact-checking is costly and does not scale easily, the allocation of time and efforts should prioritize lesser-known sources over established ones. This finding is, to our best knowledge, new to the literature on (political) misinformation (see Nyhan 2020 for a survey).
Misperception recalibration.
A large empirical literature studies the treatment effects of misperception recalibration (Bursztyn and Yang, 2022). While early research indicated the possibility of a “backfire effect” whereby recalibration efforts (e.g., provision of corrective information, cross-cutting exposure) end up reinforcing rather than reducing subjects’ misperceptions, an emerging research consensus finds that the backfire effect could be more tenuous than previously thought and that more attention should be paid to alternative issues, such as the durability and scalability of the interventions (Nyhan, 2021). We caution that even if misspecifications can be partially corrected (bring $\Delta_{i}$, $i\in\mathcal{M}$, closer to zero), their behavioral and welfare consequences may remain ambiguous due to the compound learning effect studied in Section 4.3.
5.2 Project investment with misspecified learning
A manager invests an initial wealth of $W>0$ in $K>1$ projects whose returns $\omega_{k}$, $k=1,\cdots,K$ are independently drawn from the standard normal distribution. Prior to decision making, he inquires about the return of each project from a mix of familiar sources — his personal source, internal consultants — and unfamiliar sources — external consultants, robo-advisors. The signal generated by each source is the same as in the baseline model, with its bias now capturing how optimistic the source is about the return of the project.
[…]
Over time, the manager observes signal realizations and updates his belief about source biases. His long-run belief underestimates the biases of sources that are optimistic about the returns of projects, while overestimating the biases of those that are neutral or pessimistic. We take this belief as given and examine its consequences for the manager’s investment decisions and welfare. For ease of notation, we assume, in the remainder of this section, that only sources that report on project $k=1$ have their biases misspecified by the manager. In addition, the manager has a CARA Bernoulli utility function with an Arrow-Pratt measure $\lambda>0$ of absolute risk aversion.
Specifically, for each project $k=1,\cdots,K$, we use $\mu_{k}^{*}$ to denote a typical realization of the best linear predictor of its return under the correctly specified model. The counterpart of this quantity under the misspecified model is denoted by $\widehat{\mu}_{k}$. The variance $\eta_{k}^{-1}$ of the best linear predictor is known to decrease with the signal-to-noise ratios of the signals that report on project $k$. For any given profile of $\mu_{k}$, $k=1,\cdots,K$, the optimal investment decisions, denoted by $\alpha_{k}$, $k=1,\cdots,K$, solve for
Let $(\alpha_{k}^{*})_{k=1}^{K}$ and $(\widehat{\alpha}_{k})_{k=1}^{K}$ denote the optimal portfolio choices under the correctly specified model and the misspecified model, respectively. The difference $\Delta_{\alpha_{k}}\coloneqq\widehat{\alpha}_{k}-\alpha_{k}^{*}$ captures the distortionary effect of misspecified learning on the investment in project $k$. The expected utilities generated by $(\alpha_{k}^{*})_{k=1}^{K}$ and $(\widehat{\alpha}_{k})_{k=1}^{K}$ are denoted by $\text{EU}^{*}$ and $\widehat{\text{EU}}$, respectively. Their difference captures the distortionary effect of misspecified learning on the manager’s welfare.
Proposition 5.
[…]
and their absolute values are all decreasing in $\eta_{k}^{-1}$, $k=1,\cdots,K$.
We perform four comparative static exercises. Consider first the effect of expanding the set $\mathcal{S}-\mathcal{M}$ of unfamiliar sources, due to the advent of robo-advisors or the popular advice that projects should be evaluated by “uninvolved managers” to counteract entrenched biases in managerial decision making (Guenzel and Malmendier, 2020). This change has no impact on $\overline{\Delta}$ by DOOM while clearly reducing the errors associated with the best linear predictors (as captured by $\eta_{k}^{-1}$, $k=1,\cdots,K$). Consequently, it mitigates the distortionary effects of misspecified learning on the manager’s decision making and expected utility.
Consider next the effect of increasing the signal-to-noise ratios of the sources, which causes $\eta_{k}^{-1}$, $k=1,\cdots,K$ to decrease. Although this change unambiguously mitigates the distortions of the manager’s decisions and expected utility, holding $\overline{\Delta}$ fixed, its impact on $\overline{\Delta}$ is generally ambiguous. This suggests that technological advances that lower data collection and processing costs (e.g., the AI revolution) could have differential impacts on our manager, depending on whether they operate by increasing the signal-to-noise ratios of familiar sources or those of novel sources.
Consider next corrective measures aimed at bringing misspecifications closer to zero, which in reality range from debiasing board members who are prone to select and retain like-minded CEOs (Guenzel and Malmendier, 2020) to nudges such as premortem, i.e., imagine all twists and turns (Cohee and Barnhart, 2023). To the extent that the comparative statics of $\overline{\Delta}$ against these measures are generally ambiguous in the presence of compound misspecified learning, their impacts on the manager’s decision making and welfare can be rather nuanced.
Lastly, consider the effect of expanding the portfolio of the projects overseen by the manager. Since this change mechanically increases $\sum_{k=1}^{K}\eta_{k}$, it exacerbates the distortions of the manager’s decisions and expected utility when the distributions of the signals that report on the additional projects are correctly specified. In the presence of misspecified learning, adding projects with independent returns to the portfolio creates the illusion that the optimal portfolio is well diversified. Consequently, more funds are allocated to projects the manager is overconfident about, thereby exacerbating the distortions of his decisions and welfare.
6 Concluding remarks
This paper takes the first step toward understanding the outcomes of Bayesian learning in the presence of multiple misspecified information sources. We propose a framework that encompasses a broad set of economic phenomena and develop a methodology that delivers tractable results and new insights. Several questions remain open.
The first question concerns the utilization of current findings as pivotal steps in approaching new inquiries. For example, how do managerial misperceptions interact with minority employees’ decisions to undertake productivity-enhancing investment? How does the polarization of voters’ behavior influence politicians’ choices of platforms and policies? How are equilibrium asset prices affected by distortions in individual investors’ portfolio choices? Successfully addressing these questions requires that we incorporate misspecified active learning into the current framework and tackle additional theoretical challenges such as the convergence of equilibria.
As discussed in Section 1.1, our results on labor market discrimination are distinctive from those due to tastes or rational beliefs, prompting the need for experimental validation. Some of the experimental designs proposed in Section 1.1, such as testing the comparative statics regarding informational interventions, have already been implemented by Bohren et al. (2023), albeit in a simpler decision environment devoid of misspecified learning about unfamiliar sources. Testing models of misspecified learning can be challenging, although successful endeavors have been undertaken previously; see, e.g., Bohren et al. (2019) for using the dynamics of beliefs to identify the source of discrimination. Considering the state of the literature, an empirical evaluation of our theory by means of experiments seems both interesting and feasible.
Outside the realm of misspecified Bayesian learning, misperceptions are often explained by learning frictions such as non-Bayesian updating or errors in reasoning. Ortoleva (2022) provides an up-to-date survey of these models and their economic applications. In contrast, our DM has a misspecified prior but is otherwise a standard Bayesian. We view the two approaches as complementary. Their connection is explored by Bohren and Hauser (2023), who provide conditions when a non-Bayesian updating rule has a misspecified-learning representation.
Appendix A Proofs
A.1 Proof outline
We begin by describing the problem studied by HKS and discuss the relations between our analyses. HKS examine a problem of misspecified Bayesian learning in which the DM observes in each period $t=1,2,\cdots$ a $D$-dimensional Gaussian signal with mean $\Gamma f\in\mathbb{R}^{D}$ and covariance matrix $\Sigma\in\mathbb{R}^{D\times D}$. Initially, the DM knows the transformation matrix $\Gamma\in\mathbb{R}^{D\times N}$ while misspecifying the first entry of the fundamental $f\in\mathbb{R}^{N}$. He may or may not know the true covariance matrix $\Sigma^{*}$, depending on whether he is in Case (I) or Case (III). [^14] Case (II) of HKS captures situations in which the DM holds a degenerate prior belief about the fundamental value. The support of his prior belief is
$\{(f,\Sigma)\in\mathbb{R}^{N}\times\mathbb{R}^{D\times D}:f_{1}=\tilde{f}_{1},\Sigma=\Sigma^{*}\}$
in Case (I). In Case (III), it becomes
$\{(f,\Sigma)\in\mathbb{R}^{N}\times\mathbb{R}^{D\times D}:f_{1}=\tilde{f}_{1},\Sigma\text{ is SPD with }\lambda_{\min}(\Sigma)\geq\underline{\underline{\lambda}}\},$
where $\underline{\underline{\lambda}}$ can be any positive number smaller than $1\wedge\lambda_{\min}(\Sigma^{*})$. Over time, the DM updates his belief about $(f,\Sigma)$ in a Bayesian manner based on the prior and signal realizations. For both cases, HKS demonstrate that the DM’s limiting posterior belief concentrates on the KL minimizer and provide characterizations for the latter.
The current analysis differs from that of HKS in the following aspects.
  1. D1
    We allow the DM to misspecify multiple entries of the fundamental, while HKS focus on the case of a single misspecified entry.
  2. D2
    We are interested in the exercise of predicting the true state of the world based on signal realizations. To this end, we assume that signals equal biases plus the true state, plus state-independent errors with a covariance matrix $\Omega$.
  3. D3
    $\Omega$ is taken to be arbitrary when we solve the KL minimizer and prove that it is the limit point of the DM’s posterior belief. When applying our results to derive economic insights in more specific applications, we restrict $\Omega$ to be low-rank updates of diagonal matrices.
How do these differences impact the analysis? For establishing the convergence of the limiting posterior belief to the KL minimizer, it turns out that D1 does not matter, whereas the differences due to D2 are quite nuanced. We highlight them in Appendix A.2 when we replicate the convergence proof of HKS line by line.
Our solution to the KL minimization problem nests that of HKS due to D1. Our method exploits the first-order conditions associated with KL minimization, together with the block matrix inversion formula. Its difference from HKS’ guess-and-verify approach is underscored in Appendix A.3 when we present the proof of Lemma 2.
The restriction of $\Omega$ to low-rank updates of diagonal matrices is suitable for the economic applications of our interest. The proofs presented in Appendix A.4 exploit existing mathematical results on low-rank updates of invertible matrices.
We follow the notational conventions developed in the main text. Specifically, $x$ denotes the $N$-vector $[x_{1},\cdots,x_{N}]^{\top}$, $\text{diag}(\cdot)$ denotes the diagonal matrix created from a vector, $\mathbf{1}_{k}$ and $\mathbf{0}_{k}$ denote $k$-vectors of ones and zeros, respectively, for each $k\in\mathbb{N}$, and $\mathbb{I}_{k\times k}$ denotes the $k\times k$ identity matrix. $\otimes$ denotes the Kronecker product. $\max\{a,b\}$ is written as $a\vee b$; $\min\{a,b\}$ as $a\wedge b$.
A.2 Proofs of the concentration of limiting posterior belief on KL minimizer
This appendix gathers proofs for the concentration of the DM’s limiting belief on the set of KL minimizers. We begin by presenting a useful lemma.
Lemma 5.
Let $G\in\mathbb{R}^{N\times N}$ be any symmetric matrix and $\lambda_{k}(G)$ denote the $k^{th}$ largest eigenvalue of $G$, $k=1,\cdots,N$. The following are true for any $x\in\mathbb{R}^{N}$:
@@ A.2 Proofs of the concentration of limiting posterior belief on KL minimizer
Weyl’s inequality (Weyl, 1912) states that $\lambda_{i+j-1}(A+B)\leq\lambda_{i}(A)+\lambda_{j}(B)\leq\lambda_{i+j-N}(A+B)$ for any Hermitian matrices $A$ and $B$ that belong to an inner product space of dimension $N$. Letting $A=G$ and $B=xx^{\top}$ in Weyl’s inequality yields
@@
where the equalities use the fact that $xx^{\top}$ has rank one and so has two eigenvalues: 00 and $\text{tr}(xx^{\top})$, together with $\text{tr}(xx^{\top})=\text{tr}(x^{\top}x)=x^{\top}x>0$. Interchanging the roles of $G+xx^{\top}$ and $G$ in the above derivation gives us the following:
@@
that is the first inequality in the statement of the lemma. (4), (5), and (6) imply that
@@
that is the second inequality in the statement of the lemma. ∎
@@
In the environment considered in 4.3, the support of the DM’s prior belief is
@@
where $\Lambda$ denotes the set of $N\times N$ SPD matrices whose minimum eigenvalues are bounded below by $\underline{\lambda}\in(0,1\wedge\lambda_{\min}(\Omega^{*}))$. The proof presented below verifies that this environment satisfies the assumptions of Berk (1966), hence the DM’s limiting posterior belief concentrates on the set of KL minimizers.
@@
is continuous in $(b,\Sigma)$ on $\Theta$. This is clearly true, since $\text{det}\Sigma$ and $\Sigma^{-1}$ are both continuous in $\Sigma$.
@@
Part (iii) of Berk’s assumption says for every $(b,\Sigma)\in\Theta$, there exists an open neighborhood $U\subset\Theta$ of $(b,\Sigma)$ such that
@@
Let $\lambda_{\min}(\Sigma^{\prime})$ and $\lambda_{\max}(\Sigma^{\prime})$ denote the minimum eigenvalue and the maximum eigenvalue of $\Sigma^{\prime}$, respectively. Then
$|\log l(x\mid b^{\prime},\Sigma^{\prime})|$ $=|\frac{1}{2}\log((2\pi)^{N}\text{det}\Sigma^{\prime})+\frac{1}{2}(x-b^{\prime})^{\top}(\Sigma^{\prime})^{-1}(x-b^{\prime})|$
@@
where the inequality uses the facts that the determinant of a matrix equals the product of its eigenvalues, the Rayleigh quotient $(x-b^{\prime})^{\top}(\Sigma^{\prime})^{-1}(x-b^{\prime})/(x-b^{\prime})^{\top}(x-b^{\prime})$ is bounded above by $\lambda_{\max}(\Sigma^{\prime-1})=1/\lambda_{\min}(\Sigma^{\prime})$, and $\lambda_{\min}(\Sigma^{\prime})\geq\underline{\lambda}>0$. Since $\lambda_{\min}(\Sigma^{\prime})$ and $\lambda_{\max}(\Sigma^{\prime})$ are continuous in the entries of $\Sigma^{\prime}$, the last expression is finite on every finite neighborhood $U$.
The proof so far is the same as that of HKS’ Theorem A.1 Case (III). The last step of the proof differs slightly due to our restriction of the covariance matrix to those of the form $\Omega+\mathbf{1}_{N}\mathbf{1}_{N}^{\top}$. Part (iv) of Berk’s assumption states that for every $c\in\mathbb{R}$, there exists $D\subset\Theta$ with a compact complement $\Theta-D$ such that
@@
Fix any $k_{1},k_{2}>0$, and let $D$ be the set of $(b^{\prime},\Sigma^{\prime})$ such that $\|b^{\prime}-b^{*}\|>k_{1}$ and the corresponding $\Omega^{\prime}$ satisfies $\lambda_{\max}(\Omega^{\prime})>k_{2}$. By Lemma 5, $\lambda_{\max}(\Sigma^{\prime})\geq\lambda_{\max}(\Omega^{\prime})(>k_{2})$ and $\lambda_{\min}(\Sigma^{\prime})\geq\lambda_{\min}(\Omega^{\prime})(\geq\underline{\lambda})$. Substituting these bounds into HKS’ proof of their Theorem A.1 Case (III) and replicating line by line, we obtain the following:
@@
$\leq$ $-\frac{1}{2}\left(N\log(2\pi)+(N-1)\log\lambda_{\min}(\Sigma^{\prime})+\log(k_{2}\vee\|x-b^{\prime}\|^{2})\right),$
where the first inequality uses the facts that $\text{det}\Sigma^{\prime}\geq\lambda_{\min}^{N-1}(\Sigma^{\prime})\lambda_{\max}(\Sigma^{\prime})$, $\lambda_{\min}(\Sigma^{\prime})\geq\underline{\lambda}>0$, and the Rayleigh quotient $(x-b^{\prime})^{\top}(\Sigma^{\prime})^{-1}(x-b^{\prime})/(x-b^{\prime})^{\top}(x-b^{\prime})$ is bounded below by $\lambda_{\min}(\Sigma^{\prime-1})=1/\lambda_{\max}(\Sigma^{\prime})$. The second inequality follows from minimizing $\log\lambda_{\max}(\Sigma^{\prime})+\|x-b^{\prime}\|^{2}/\lambda_{\max}(\Sigma^{\prime})$ over the range $\lambda_{\max}(\Sigma^{\prime})>k_{2}$. To obtain a lower bound for $\|x-b^{\prime}\|^{2}$, we note that $x-b^{\prime}=x-b^{*}+(b^{*}-b^{\prime})=(\omega+\varepsilon)+(b^{*}-b^{\prime})$ and so $\|x-b^{\prime}\|^{2}\geq\|\omega+\varepsilon\|^{2}+\|b^{*}-b^{\prime}\|^{2}-2\|\omega+\varepsilon\|\|b^{*}-b^{\prime}\|$ by the Cauchy-Schwarz inequality. Since $\|b^{*}-b^{\prime}\|>k_{1}$ by assumption, it follows that for any given value of $\|\omega+\varepsilon\|<k_{1}/2$, $\|b^{*}-b^{\prime}\|^{2}+\|\omega+\varepsilon\|^{2}-2\|b^{*}-b^{\prime}\|\|\omega+\varepsilon\|$ is minimized at $\|b^{*}-b^{\prime}\|=k_{1}$ and so is bounded below by $k_{1}^{2}/4$. Consequently, the last line of the above inequality is bounded above by

### v3

@@ A.2 Useful lemmas and their proofs
Fix any symmetric matrix $G\in\mathbb{R}^{N\times N}$ and let $\lambda_{k}(G)$ denote the $k^{th}$ largest eigenvalue of $G$, $k=1,\cdots,N$. The following statements are true for any $x\in\mathbb{R}^{N}$:
@@
Weyl’s inequality (Weyl, 1912) states that $\lambda_{i+j-1}(A+B)\leq\lambda_{i}(A)+\lambda_{j}(B)\leq\lambda_{i+j-N}(A+B)$ for any Hermitian matrices $A$ and $B$ belonging to an inner product space of dimension $N$. Letting $A=G$ and $B=xx^{\top}$ yields
@@
where the equalities exploit the fact that $xx^{\top}$ has rank one and so has two eigenvalues: 00 and $\text{tr}(xx^{\top})$, and that $\text{tr}(xx^{\top})=\text{tr}(x^{\top}x)=x^{\top}x>0$. Interchanging the roles of $G+xx^{\top}$ and $G$ in the above derivation yields
@@
that is, the first inequality in the statement of the lemma. Additionally, (13), (14), and (15) imply that
@@
that is, the second inequality in the statement of the lemma. ∎
@@
In the environment considered in Section 5.1, the support of the DM’s prior beliefs is
@@
where $\Lambda$ denotes the set of $N\times N$ SPD matrices whose minimum eigenvalues are bounded below by $\underline{\lambda}\in(0,1\wedge\lambda_{\min}(\Omega^{*}))$. The proof presented below verifies that this environment satisfies the assumptions of Berk (1966), hence the DM’s limiting posterior beliefs concentrate on the set of KL minimizers.
@@
is continuous in $(b,\Sigma)$ on $\Theta$. This is clearly true in our setting, since $\text{det}\Sigma$ and $\Sigma^{-1}$ are both continuous in $\Sigma$.
@@
Part (iii) of Berk’s assumption says that for every $(b,\Sigma)\in\Theta$, there exists an open neighborhood $U\subset\Theta$ such that
@@
Let $\lambda_{\min}(\Sigma^{\prime})$ and $\lambda_{\max}(\Sigma^{\prime})$ denote the minimum and maximum eigenvalues of $\Sigma^{\prime}$, respectively. Then
$|\log l(x\mid b^{\prime},\Sigma^{\prime})|$ $=|\frac{1}{2}\log((2\pi)^{N}\text{det}\Sigma^{\prime})+\frac{1}{2}(x-b^{\prime})^{\top}\Sigma^{{}^{\prime}-1}(x-b^{\prime})|$
@@
where the inequality uses the following facts: (i) the determinant of a matrix equals the product of its eigenvalues; (ii) the Rayleigh quotient $(x-b^{\prime})^{\top}\Sigma^{{}^{\prime}-1}(x-b^{\prime})/(x-b^{\prime})^{\top}(x-b^{\prime})$ is bounded above by $\lambda_{\max}(\Sigma^{\prime-1})=1/\lambda_{\min}(\Sigma^{\prime})$; (iii) $\Sigma^{\prime}$ is SPD, and so $\lambda_{\min}(\Sigma^{\prime})>0$. Since $\lambda_{\min}(\Sigma^{\prime})$ and $\lambda_{\max}(\Sigma^{\prime})$ are continuous in the entries of $\Sigma^{\prime}$, the last expression is finite on every finite neighborhood $U$.
So far, the proof follows the same steps as HKS’ proof for their Theorem A.1 Case (III). The last step of the proof differs slightly, due to our restriction of the covariance matrix to those of the form $\Omega+\mathbf{1}_{N}\mathbf{1}_{N}^{\top}$. Part (iv) of Berk’s assumption says that for every $c\in\mathbb{R}$, there exists $D\subset\Theta$ with a compact complement $\Theta-D$ such that
@@
Fix any $k_{1},k_{2}>0$. Let $D$ be the set of $(b^{\prime},\Sigma^{\prime})$ such that $\|b^{\prime}-b^{*}\|>k_{1}$ and the corresponding $\Omega^{\prime}$ satisfies $\lambda_{\max}(\Omega^{\prime})>k_{2}$. By Lemma 3, $\lambda_{\max}(\Sigma^{\prime})\geq\lambda_{\max}(\Omega^{\prime})>k_{2}$, and $\lambda_{\min}(\Sigma^{\prime})\geq\lambda_{\min}(\Omega^{\prime})\geq\underline{\lambda}$. Substituting these bounds into HKS’ proof and replicating line by line, we obtain that
@@
$\leq$ $-\frac{1}{2}\left(N\log(2\pi)+(N-1)\log\lambda_{\min}(\Sigma^{\prime})+\log(k_{2}\vee\|x-b^{\prime}\|^{2})\right).$
In the above derivation, the first inequality uses the following facts: (i) $\text{det}\Sigma^{\prime}\geq\lambda_{\min}^{N-1}(\Sigma^{\prime})\lambda_{\max}(\Sigma^{\prime})$; (ii) $\lambda_{\min}(\Sigma^{\prime})\geq\underline{\lambda}>0$; and (iii) the Rayleigh quotient $(x-b^{\prime})^{\top}(\Sigma^{\prime})^{-1}(x-b^{\prime})/(x-b^{\prime})^{\top}(x-b^{\prime})$ is bounded below by $\lambda_{\min}(\Sigma^{\prime-1})=1/\lambda_{\max}(\Sigma^{\prime})$. The second inequality follows from minimizing $\log\lambda_{\max}(\Sigma^{\prime})+\|x-b^{\prime}\|^{2}/\lambda_{\max}(\Sigma^{\prime})$ over the range $\lambda_{\max}(\Sigma^{\prime})>k_{2}$. To obtain a lower bound for $\|x-b^{\prime}\|^{2}$, we note that $x-b^{\prime}=x-b^{*}+(b^{*}-b^{\prime})=(\omega+\epsilon)+(b^{*}-b^{\prime})$, hence $\|x-b^{\prime}\|^{2}\geq\|\omega+\epsilon\|^{2}+\|b^{*}-b^{\prime}\|^{2}-2\|\omega+\epsilon\|\|b^{*}-b^{\prime}\|$ by the Cauchy-Schwarz inequality. Since $\|b^{*}-b^{\prime}\|>k_{1}$ by assumption, for any given value of $\|\omega+\epsilon\|<k_{1}/2$, $\|b^{*}-b^{\prime}\|^{2}+\|\omega+\epsilon\|^{2}-2\|\omega+\epsilon\|\|b^{*}-b^{\prime}\|$ is minimized at $\|b^{*}-b^{\prime}\|=k_{1}$ and so is bounded below by $k_{1}^{2}/4$. Consequently, the last line of the above inequality is bounded above by